The ten's digit of a two digit number is 2 less than the unit's digit. The value of the number is 7 times the unit's digit. find the digits and the value of the number.
Answer by stanbon(75887)

The ten's digit of a two digit number is 2 less than the unit's digit. The value of the number is 7 times the unit's digit. find the digits and the value of the number.
------------
Let the units digit be "x".
The ten's digit is "x-2".
The value of the number is 10(x-2)+x=11x-20
EQUATION: 11x-20=7x
4x=20
x=5 (this is the units digit)
x-2=3 (this is the 10's digit)
The value of the number is 35
-----------
Cheers,
Stan H.